What`s the electric field intensity between the plates with separation of 2 cm of each 150 V and -50 V respectivlry?

Grade:12

1 Answers

DR STRANGE
7 years ago
Hi Ijas, 
 you can use V=Ed i.e. potential diff between the plates is distance times the electric field between them.
V=150-(-50)=200, d=2cm = 0.02m , therefore , 200= E*0.02   =>E=104V/m
please approve if you are satisfied :)